a piston-cylinder device initially contains a rate 150 kilo pascal and 27 degree celsius at this state the volume is 400 later the mass of the piston is such that the 50 kilo pascal pressure is required to move the area is nau heated until its volume is doubled determine the total heat transfer to the air
1
Expert's answer
2020-04-15T01:52:26-0400

Doubled volume equals 800

PV = nRT

n = PV/RT

Assume the volume is 0.4 L!

n = 150000*0.4x10-3 / 8.314*300 = 0.024 mol of air

M = 29

m = nM = 0.024*29 = 0.698 g of air.

T2 = V2*T1/V1

T2 = 800*300/400 = 600 K

Cp for air at 300 is 1.005 kJ/kg*K

Q = Cp*m*(T2-T2)

Q = 1.005 * 0.698/1000 *(600-300)

Q = 0.21 kJ.